are Prove that Ā = 4ax – 2ay – az, B = ax + 4ay – 4az perpendicular to each other.
Answer #1

+ we show Ā = 400- zay - az B= ax + Gay - 492. To show I and R are perpendicular that A B = 0. ( dot product) (49n-zay-az) (a